OpenStudy (gabylovesyou):

Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. My birthday is June 8, 1999. How do i do dis ? ;p

OpenStudy (gabylovesyou):

@sleepyjess

OpenStudy (gabylovesyou):

@chosenmatt

OpenStudy (chosenmatt):

gimme a sec ;)

OpenStudy (gabylovesyou):

ok thank you ^-^

OpenStudy (chosenmatt):

u know....i rather get someone else too help u because i don't wanna give u false information because im not 100% sure sorry :(

OpenStudy (gabylovesyou):

its okay.. who do you think can help me ? ;p

OpenStudy (chosenmatt):

@perl @adrynicoleb @dan815 @Trinity_haha @inowalst @iGreen

OpenStudy (chosenmatt):

@js7474 please help :)

OpenStudy (chosenmatt):

@TheSmartOne please help

OpenStudy (anonymous):

Sorry hun i have no clue O.o

TheSmartOne (thesmartone):

Vertex form is \(\sf a(x-h)^2+k\)

OpenStudy (chosenmatt):

yay i got someone too help u :)

TheSmartOne (thesmartone):

The 'a' should be kept 1 as it is way easier than choosing another value between -4 and 4.

TheSmartOne (thesmartone):

'k' is 8 which is the day in your birthday.

TheSmartOne (thesmartone):

And 'h' is 6 because June is the 6th month.

OpenStudy (gabylovesyou):

@chosenmatt Thank you so much !

OpenStudy (gabylovesyou):

@TheSmartOne sooo it would be f(x) = 1(x - 6)^2 + 8

OpenStudy (gabylovesyou):

right ?

TheSmartOne (thesmartone):

Correct :)

OpenStudy (gabylovesyou):

ok... how do i turn that equation in standard form ?

TheSmartOne (thesmartone):

Standard Form is \(\sf Ax^2+By=C\)

TheSmartOne (thesmartone):

So expand \(\sf (x-6)^2\) and combine it with the 8 and you can also think of \(\sf f(x) = y\) and get it into that format above.

OpenStudy (chosenmatt):

:-)

OpenStudy (gabylovesyou):

(x - 6)(x - 6) ?

OpenStudy (gabylovesyou):

I get x^2 - 12x + 44

OpenStudy (gabylovesyou):

@TheSmartOne

OpenStudy (gabylovesyou):

so y = x^2 - 12x + 44

TheSmartOne (thesmartone):

Now move around the terms so it is in the format of \(\sf Ax^2+By=C\)

OpenStudy (gabylovesyou):

umm.. x^2 + 44 - 12x = y ?

TheSmartOne (thesmartone):

Correct. Ignore what I was saying lol. Good job :) What is the next part?

OpenStudy (gabylovesyou):

so it is like that ? ok lol ;p how do i graph it now.

TheSmartOne (thesmartone):

https://www.desmos.com/calculator/ju5tzfaf2d

TheSmartOne (thesmartone):

@iambatman explains it nicely over here too: http://openstudy.com/study#/updates/53af97e9e4b02c45b016004b

OpenStudy (gabylovesyou):

Thank you so much :D

TheSmartOne (thesmartone):

Anytime :)